Step 2 Since 29 is two standard deviations to the left of μ and 41 is two standard deviations to the right on, we need to find the area under the normal curve from μ- ơ) to +20 Area Under a Normal Curve 2.35%/ 13.5% | 34% | 34% | 13,5% 2.35% 68% 95% 99.7% X % of the area under the curve falls between these values. Therefore, days. In other words, we expect 1 X % of the 10,000 chicks can be expected to hatch in 29 to 41 x eggs to hatch during this time period.
Answer #1

Answer : 95, 95, 9500

% of area that lie between 29 and 41 = 13.5 + 34 + 34 + 13.5 = 95% area lies between 2 standard deviation from mean.

95 % of 10000 chicks are expected to hatch between 29 and 41 days. 95/100 * 10000 = 9500 chicks